OpenStudy (anonymous):

write an explicit formula for the sequence 8, 6, 4, 2, 0, ... Then find a(down)14

jimthompson5910 (jim_thompson5910):

you're subtracting by 2 each time

jimthompson5910 (jim_thompson5910):

so the common difference is d = -2

jimthompson5910 (jim_thompson5910):

the first term is 8, so a1 = 8

jimthompson5910 (jim_thompson5910):

an = a1 + d(n-1) an = 8 + (-2)(n-1) an = 8 - 2n + 2 an = -2n + 10 So the explicit formula is an = -2n + 10 Does this make sense?